Among coffee drinkers, men drink a mean of 3.2 cups per day with a standard deviation of 0.8 cups. Assume the number of drinking per day follows a normal distribution.

If the top 5% of coffee drinkers are considered heavy coffee drinkers, what is the minimum number of cups consumed by a heavy coffee drinker?
Answer #1

Solution :

Given that,  

mean = = 3.2

standard deviation = = 0.8

Using standard normal table ,

P(Z > z) = 5%

1 - P(Z < z) = 0.05

P(Z < z) = 1 - 0.05 = 0.95

P(Z < 1.65) = 0.95

z = 1.65

Using z-score formula,

x = z * +

x = 1.65* 0.8 + 3.2 = 4.52

The minimum number of cups consumed by a heavy coffee drinker is 4.52